OpenStudy (anonymous):

three consecutive even integers have a sum of at most 180. what is the largest possible integer?

OpenStudy (anonymous):

set it up as (x) + (x+1) + (x+2) = 180 and then sove for x and plug in to get three cinsecutive answers

myininaya (myininaya):

let n be the first even integer so n+2 is the next even integer n+4 is the next next even integer

myininaya (myininaya):

n+(n+2)+(n+4)=180 3n+6=180 solve for n and then find n+4

OpenStudy (anonymous):

Thank you so much, the other person didnt make the integers even since 1 is odd, but you did thank you!!!!

myininaya (myininaya):

you are right fresh he has even+odd+even=180

OpenStudy (anonymous):

Solved in one statement with Mathematica:\[\text{Solve}[\text{Plus}[x-4,x-2,x]=180,x]\]\[\{x\to 62\} \]